The force F = 50 N acts on the material point. Find the change of its pulse over a time delta t = 2 s. What path has the subject point traversed during this time? Its mass is m = 600 g, while the initial velocity was v = 2m / s.  


1
Expert's answer
2021-01-31T03:39:25-0500

The acceleration of the point

"a=\frac{F}{m}=\frac{50\:\rm N}{0.6\:\rm kg}=83.3\:\rm m/s^2"

The distance traveled by point

"d=v_0t+\frac{at^2}{2}=2*2+\frac{83.3*2^2}{2}=171\:\rm m"